# K3 surfaces associated to a cubic fourfold

### v1

@@ 5.1. The case p=3
Let $S\in\mathcal{F}_{g}$, where $g=n^{2}+n+2$ with a symplectic automorphism $\sigma$ of order 3 and let $X\in\mathcal{C}_{d}$ be the image of $S$ under the rational map in 5.1. Then $F(X)\simeq S^{[2]}$ and $\sigma$ induces a natural automorphism $\sigma^{[2]}$ on $F(X)$ with 27 isolated fixed points. The automorphism $\sigma^{[2]}$ is induced by an automorphism $\tau$ of $X$, which in turn is induced by a linear automorphism of $\mathbf{P}^{5}$.Then one can choose coordinates $[x_{0},x_{1},x_{2},x_{3},x_{4},x_{5}]$ in such a way that $(X,\tau)$ belongs to one of the families of cubic fourfolds with a symplectic automorphism of order 3, as described in [LFu, Thm.1.1]. There are two families of cubic fourfolds with a symplectic automorphism of order 3, whose fixed locus on $F(X)$ consists of 27 isolated fixed points.
@@
with $f$ of degree 3 and $L$ of degree 1. The fixed locus of the automorphism $\tau_{1}$ on $X$ is the cubic surface $S:f(x_{0},x_{1},x_{2},x_{3})=0$. Since this fixed locus is not finite the automorphism induced by 5.10 on $F(X)$ cannot coincide with $\sigma^{[2]}$ of $F(X)$,i.e. it is not natural. In [Mong,7.2.8] it is proved that the order 3 symplectic automorphism of the cubics in $F_{3}$ are standard.
@@
[…]
Since the moduli space of polarized K3 surfaces with a symplectic automorphism of order 3 has dimension 7(see [GS],Cor.5.1]), the image of the the map $\mathcal{F}_{g}\to G_{3}$ has codimension 1.
@@
which is defined outside the 27 fixed points on $S^{[2]}$.The rational map $\pi^{[2]}$ induces, via the isomorphisms in 5.12, a rational map $F(\tilde{X})\dashrightarrow F(X)$.
Since the transcendental motive $t_{2}(-)$ of a surface is a birational invariant the maps $\tilde{S}\to S$ and $\tilde{S}\to Y$ in 5.9 induce a map $\theta:t_{2}(S)\to t_{2}(Y)$ that is the projection onto a direct summand, see [Ped,Prop.1]. Therefore

### v2

@@ 5.1. The case p=3
Let $S\in\mathcal{F}_{g}$, where $g=n^{2}+n+2$ with a symplectic automorphism $\sigma$ of order 3 and let $X\in\mathcal{C}_{d}$ be the image of $S$ under the rational map in 5.1. Then $F(X)\simeq S^{[2]}$ and $\sigma$ induces a natural automorphism $\sigma^{[2]}$ on $F(X)$ with 27 isolated fixed points. The automorphism $\sigma^{[2]}$ is induced by an automorphism $\tau$ of $X$, which in turn is induced by a linear automorphism of $\mathbf{P}^{5}$. Then one can choose coordinates $[x_{0},x_{1},x_{2},x_{3},x_{4},x_{5}]$ in such a way that $(X,\tau)$ belongs to one of the families of cubic fourfolds with a symplectic automorphism of order 3, as described in [Fu, Thm.1.1]. There are two families of cubic fourfolds with a symplectic automorphism of order 3, whose fixed locus on $F(X)$ consists of 27 isolated fixed points.
@@
with $f$ of degree 3 and $L$ of degree 1. The fixed locus of the automorphism $\tau_{1}$ on $X$ is the cubic surface $S:f(x_{0},x_{1},x_{2},x_{3})=0$. Since this fixed locus is not finite the automorphism induced by 5.11 on $F(X)$ cannot coincide with $\sigma^{[2]}$ of $F(X)$, hence it is not natural. In [Mong,7.2.8] it is proved that the order 3 symplectic automorphism of the cubics in $F_{3}$ are standard.
@@
where $f_{1},f_{2},f_{3}$ have degree 3 and $i=0,1;j=2,3;k=4,5$. The fixed locus of $\tau_{2}$ on $X$ is finite and consists of the 9 isolated points $P_{i,1},P_{i,2},P_{i,3}$ satisfying the the equation $f_{i}=0$, for $i=1,2,3$ The fixed locus on $F(X)$ corresponds to the 27 lines $\overline{P_{i,k}P_{j,l}}$, for $1\leq i<j\leq 3$ and $k,l=1,2,3$, see [Fu,IV-(5)]. Each point on $X$ fixed by the automorphism is the intersection of 3 lines in $F(X)$ fixed by the automorphism.
Since the moduli space of polarized K3 surfaces with a symplectic automorphism of order 3 has dimension 7 (see [GS],Cor.5.1]), the image in $G_{3}$ has codimension 1 and contains the cubics $X$ such that the automorphism $\tau_{2}$ is natural.
@@
which is defined outside the 27 fixed points on $S^{[2]}$. The rational map $\pi^{[2]}$ induces, via the isomorphisms in 5.13, a rational map $F(\tilde{X})\dashrightarrow F(X)$.
Since the transcendental motive $t_{2}(-)$ of a surface is a birational invariant the maps $\tilde{S}\to S$ and $\tilde{S}\to Y$ in 5.10 induce a map $\theta:t_{2}(S)\to t_{2}(Y)$ that is the projection onto a direct summand, see [Ped,Prop.1]. Therefore